OpenStudy (pottersheep):

Grade 12 Vectors . Find a point P on the line L situated 19√2 units distance from plane W.? Find a point P on the line L situated 19√2 units distance from plane W. W = 5x + 3y - 4z = 18 L = (x-1)/2 = (y+2)/-1 = z/-3 Please help me. How do I do this? Thank you so much

OpenStudy (amistre64):

find the point on the plane that the xyz of the line fits

OpenStudy (pottersheep):

For paramatric equation of the line I got X = 1+2t Y = -2 –t Z = -3t

OpenStudy (amistre64):

the normal of the plane is <5,3,-4> a line with that normal is X = 1+5t Y = -2 +3t Z = 0 -4t would you agree?

OpenStudy (pottersheep):

yes

OpenStudy (amistre64):